### v1

# Federated Unlearning Made Practical: Seamless Integration via Negated Pseudo-Gradients

Abstract
The right to be forgotten is a fundamental principle of privacy-preserving regulations and extends to Machine Learning (ML) paradigms such as Federated Learning (FL). While FL enhances privacy by enabling collaborative model training without sharing private data, trained models still retain the influence of training data. Federated Unlearning (FU) methods recently proposed often rely on impractical assumptions for real-world FL deployments, such as storing client update histories or requiring access to a publicly available dataset. To address these constraints, this paper introduces a novel method that leverages negated Pseudo-gradients Updates for Federated Unlearning (PUF). Our approach only uses standard client model updates, anyway employed during regular FL rounds, and interprets them as pseudo-gradients. When a client needs to be forgotten, we apply the negated of their pseudo-gradients, appropriately scaled, to the global model. Unlike state-of-the-art mechanisms, PUF seamlessly integrates with FL workflows, incurs no additional computational and communication overhead beyond standard FL rounds, and supports concurrent unlearning requests. We extensively evaluated the proposed method on two well-known benchmark image classification datasets (CIFAR-10 and CIFAR-100) and a real-world medical imaging dataset for segmentation (ProstateMRI), using three different neural architectures: two residual networks and a vision transformer. The experimental results across various settings demonstrate that PUF achieves state-of-the-art forgetting effectiveness and recovery time, without relying on any additional assumptions, thus underscoring its practical applicability.
Index Terms:
Federated Learning, Federated Unlearning, Client Unlearning, Machine Unlearning, Privacy.
I Introduction
In today’s digital landscape, privacy has become a major concern, as reflected by the emergence of robust regulatory frameworks worldwide [1]. The European Union (EU) has consistently emphasized the importance of protecting personal data, exemplified by the introduction of the General Data Protection Regulation (GDPR) in 2016 [2]. Most recently, in May 2024, the EU enacted Regulation 2024/1183 [3], establishing the European Digital Identity Framework that empowers individuals with fine-grained control over their data. One of the key rights of these regulations is the right to be forgotten, which allows individuals to request the deletion of their previously shared data. Similar rights are central to other major privacy laws worldwide, such as the California Consumer Privacy Act (CCPA) [4] where the right to delete grants California residents the on-demand removal of personal data held by businesses.
Individuals should retain the ability to exercise their privacy rights even in Machine Learning (ML), including withdrawing their contributions from a trained model due to security or privacy concerns [5]. This can be achieved through Machine Unlearning (MU) [6, 7, 8], an emerging paradigm designed to selectively erase the influence of specific training data from a model by post-processing the trained model.
Figure 1: Performance comparison of PUF against state-of-the-art mechanisms using ResNet-18 on CIFAR-100 (non-IID case) in a 10-client setup, where one client requests unlearning. The ideal FU algorithm should (1) minimize the difference with the gold standard retrained model, across forgetting metrics such as Forget Accuracy and various MIAs (in Figure expressed as $\Delta$ Forget Accuracy and $\Delta$ MIAs), (2) ensure a short recovery phase (Recovery Rounds) and (3) preserve the generalization performance (reported in Figure as $\Delta$ Test Accuracy just after unlearning, before recovery starts). A smaller polygon represents better unlearning performances. Experimental details and additional comparisons are available in Section V.
The right to be forgotten should also extend to privacy-preserving ML paradigms such as Federated Learning (FL) [9, 10], which enables the collaborative training of a global model across multiple clients without requiring them to share their private data with a central server. In FL, clients train local models using their on-premises data and only share updates (e.g., weight differences) [11]. Despite offering enhanced privacy, the global model inevitably reflects the influence of client training data [12].
However, the intrinsic characteristics of FL, such as its non-deterministic and iterative training process, make traditional MU techniques less effective in this context [13]. This has led to the novel concept of Federated Unlearning (FU) [13, 14], which encompasses unlearning methods tailored for FL settings. The goal of an FU algorithm depends on what we are trying to remove from the global model, which could be either some samples, all the contributions of a client, or all samples belonging to a class. In this work, we address client unlearning, where a client seeks to remove all traces of its previously shared data from the global model.
Work No Historical Information No Full Participation No Stateful Clients Task Agnostic No Auxiliary Data No Computational Overhead Simultaneous Unlearning Liu et al. [15] $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ FedEraser [16] $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ FUKD [17] $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ PGA [18] $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ NoT [19] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ PUF (ours) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
TABLE I: Comparison with other works performing client unlearning in federated settings.
An effective mechanism should achieve unlearning while maintaining the overall performance of the global model. Although some FU techniques have started to be proposed, most of them rely on hard assumptions as they usually require maintaining additional information such as historical updates [16, 15, 17] or stateful clients [18]. Moreover, these works fail to address scenarios where multiple unlearning requests occur in the same time window. This condition cannot be overlooked as clients should be allowed to remove their contributions regardless of others’ willingness.
To fill this gap, this paper presents a novel method that leverages negated Pseudo-gradients Updates for Federated Unlearning (PUF). Our approach ensures seamless integration into existing FL frameworks as it enables unlearning without modifying the traditional Federated Averaging (FedAvg) [9] training protocol. Unlike most previous works, PUF eliminates the need for storing historical information (e.g., historical updates), accessing proxy data, maintaining stateful clients, or incurring impractical computational requirements. Since PUF can be seamlessly integrated into standard FedAvg, it also explicitly addresses multiple unlearning requests that concurrently occur. The client who wants to be forgotten sends an unlearning request to the server and performs one final round of FL. The server then adjusts its model updates by flipping and scaling them according to an unlearning rate.
To demonstrate the broad applicability of our approach, we implemented PUF and extensively evaluated its performance on two benchmark image classification datasets (CIFAR-10, CIFAR-100) and a real-world medical imaging dataset (ProstateMRI) for segmentation. The evaluation involved two residual networks and a vision transformer. The reported results clearly show the efficiency and effectiveness of PUF in removing client contributions to the global model across a variety of settings. Figure 1 visually compares the performance of PUF against state-of-the-art mechanisms.
Contributions. The main contributions of this work can be summarized as follows:
  • We introduce PUF, a novel approach that enables FU without requiring any modifications or additional overhead to the original FedAvg algorithm.
  • We evaluated PUF across different federated settings and compared its performance to two state-of-the-art baselines for client unlearning. We show that PUF is more effective in removing the contribution of the target clients and more efficient in recovering the expected performance.
  • We open-sourced our code to support further related research in the community. The PUF code is available for the research community at: https://github.com/alessiomora/puf_unlearning.
Organization. The remainder of this work is organized as follows. Section III provides the background on FL and FU. Section II reviews the main related work in the field. Section IV presents PUF, while Section V evaluates its performance and Section VI concludes the paper.
II Related Work
Table I presents a comparative analysis of PUF against existing approaches in the field. The columns indicate whether each method does not rely on a given requirement ($\checkmark$) or does depend on it ($\times$). Our comparison highlights key aspects of FU methods, with a primary focus on whether an approach requires and leverages historical updates [15, 16, 17]. While a parameter server may have the capacity to store all client contributions across training rounds and FL processes, this requirement also introduces a critical limitation: the server must maintain the ability to link past contributions with specific clients, conflicting with the ephemeral nature of FL model updates [20]. Moreover, since historical updates are pre-stored, these methods require defining a specific set of data for removal beforehand, typically a target client’s data. This constraint prevents them from handling more general sample unlearning scenarios, where the data to be forgotten is not pre-determined, such as when a client requests the removal of a fraction of its dataset. Wu et al. [17] further assume access to auxiliary unlabeled data to mitigate performance degradation due to unlearning. However, this assumption is often impractical, as the public dataset must exhibit semantic similarity to clients’ private data for effective use.
Given these limitations, we focus our comparison on more realistic FU methods. Halimi et al. [18] propose Project Gradient Ascent (PGA) to erase client influence by constraining local weight updates within an L2-norm ball around a reference model. While PGA does not require historical updates or auxiliary data—making it more suitable for FL—it introduces other constraints. Clients must be stateful, storing their latest model updates, and full participation is mandatory, as unlearning relies on a reference model obtained by removing the last unlearning client’s update from the current global model. Additionally, directly applying gradient ascent on forgetting data risks gradient explosion, as the loss function often lacks an upper bound, leading to divergence and poor generalization. To counteract this, Halimi et al. employ gradient clipping and L2-norm constraints [18]. However, these techniques require careful and extensive hyperparameter tuning, such as selecting the appropriate L2-norm radius or gradient clipping threshold. PUF eliminates the risk of uncontrolled gradient growth while maintaining the standard local training procedure. This is achieved by rescaling the client’s local update (pseudo-gradient) after training, rather than modifying the training dynamics. Recently, Khalil et al. [19] proposed NoT, a novel method aiming to achieve unlearning without impractical assumptions by implementing server-side unlearning. Specifically, NoT inverts the sign of the first layer’s global model parameters upon receiving an unlearning request. While this approach does not require explicit client-side unlearning, it fails to remove a target client’s contribution selectively — NoT produces the same unlearned model regardless of which client exits the federation. Our experimental evaluation (see Figure 1 and Section V) shows that NoT requires a longer recovery period and fails to match the forgetting performance of full model retraining.
In contrast to previous methods, PUF offers several key advantages. It does not rely on historical updates or auxiliary data, does not require full client participation, and operates in a stateless manner. Moreover, PUF is task-agnostic, meaning it does not need modifications based on the specific learning task. It maintains the same computational cost as standard FedAvg rounds and requires minimal tuning, as demonstrated in our experimental results. Furthermore, PUF supports sample unlearning and can handle multiple unlearning requests simultaneously. These properties make PUF a practical and effective solution for federated unlearning, inherently aligning with the design principles of FL.
III Background
III-A Federated Learning
FedAvg [9] is a widely used baseline in FL. It adopts a client-server framework in which a central server manages the global model parameters. Training proceeds in synchronous rounds, where a randomly selected subset of client devices is invited to participate. During each round, the activated clients fine-tune the global model using their local data over a fixed number of epochs. The server then collects the locally computed updates from the clients, and incorporates them into the global model $w$, before distributing the updated global parameters again.
In FedAvg, clients typically send model updates to the server. These updates are calculated as the difference between the global weights, received at the beginning of the round, and the locally trained weights. This can be expressed mathematically as:
$w_{t+1}=\frac{1}{|S_{t}|}\sum_{i\in S_{t}}w_{t}^{i}=w_{t}+\frac{1}{|S_{t}|}\sum_{i\in S_{t}}(w_{t}^{i}-w_{t})$ (1)
where $S_{t}$ represents the set of clients selected for round $t$, $w$ represents the weight of the global model and $w^{i}$ are the weights computed locally at client $i$. For clarity, in Equation 1, model updates are shown without being scaled by the number of local data samples, but this simplification does not affect the subsequent analysis. By defining client updates as $\Delta_{t}^{i}:=w_{t}^{i}-w_{t}$ and their aggregated form as:
$\Delta_{t}:=\frac{1}{|S_{t}|}\sum_{i\in S_{t}}\Delta_{t}^{i},$ (2)
and by introducing a (global) learning rate $\eta_{s}$, the FedAvg’s update rule can be rewritten as:
$w_{t+1}=w_{t}+\eta_{s}\Delta_{t}=w_{t}-\eta_{s}(-\Delta_{t})$ (3)
This shows that the server’s update rule in FedAvg is equivalent to applying SGD to the pseudo-gradient $-\Delta_{t}$ with $\eta_{s}=1$. This perspective reveals that FedAvg is a specific instance of the broader FedOpt framework [21], which can use various server-side optimizers.
III-B Federated Unlearning
An FU method aims to remove specific learned information from the global model $w$ while preserving the good knowledge acquired on data that should not be forgotten. FU objectives are classified based on the information the algorithm is expected to forget: sample unlearning, class unlearning [22], feature unlearning [23], and client unlearning. Our work focuses on client unlearning.
During a training round $t$, a client $u$ may withdraw previous contributions from the global model $w$. This is achieved through an unlearning procedure $\mathcal{U}(w_{t},D_{u})$, which can be executed by different entities within the FL framework, such as the server, the client to be forgotten, or the remaining clients [24]. An unlearning algorithm should produce a novel version of the original global model $w_{t}^{\overline{u}}$, called the unlearned model, that effectively excludes the influence of the forget data $D_{u}$. The method is effective if $w_{t}^{\overline{u}}$ exhibits performances approximately indistinguishable from the retrained model $w_{t}^{r}$, a model trained as if client $u$ had never joined the federation.
The efficiency of FU is usually measured by the number of rounds needed to obtain the generalization performance of $w_{t}^{r}$. Effectiveness in forgetting the client data $D_{u}$ is assessed by comparing the performance of $w_{t}^{\overline{u}}$ and $w_{t}^{r}$ on $D_{u}$. Common metrics for verifying the success of unlearning include loss, accuracy, and susceptibility to membership inference attacks (MIAs) [25, 26, 27]. MIAs test whether certain samples were used at training time, by providing a measure of the unlearning effectiveness in removing traces of such samples.
III-C Terminology
In this subsection, we define key terminology that will be used throughout the paper.
  • Unlearning Client: The client that requests unlearning, also referred to as the Target Client or client $u$. The objective of an unlearning algorithm is to remove the influence of client $u$ data (forgetting data) from the global model.
  • Original Model: The global model trained with the participation of both the unlearning client and all other clients.
  • Retrained Model: The global model trained without the unlearning client from the beginning of the FL process. This serves as the gold-standard baseline for evaluating the effectiveness of unlearning methods.
  • Unlearned Model: The original model after the unlearning procedure (for example, after applying PUF). Since unlearning may impact model performance, a recovery phase is often necessary to restore its generalization capabilities to match the retrained model. As detailed in Section V, an optimal FU algorithm should minimize the discrepancy with the retrained model in terms of forgetting metrics (e.g., accuracy on client $u$’s data, susceptibility to MIAs) while achieving equal or superior generalization to the retrained model in the fewest possible FL rounds.
IV PUF: Federated Unlearning via Negated Pseudo-Gradients
IV-A Preliminaries
Before presenting our original method, we provide the necessary preliminaries and notation to fully and more easily understand our proposal. During a given round $t$, a subset of the participating clients may request the removal of their contributions. We define $S_{t}^{-}$ as the subset of clients requesting unlearning, and $S_{t}^{+}$ as the remaining clients. Additionally, let $n$ denote the total number of samples held by participating clients during the unlearning round, such that:
$n:=\sum_{i\in S_{t}^{+}}|D_{i}|+\sum_{j\in S_{t}^{-}}|D_{j}|$ (4)
We define two aggregated model updates: $\Delta_{t}^{+}$ , representing the aggregated updates contributed by the clients in $S_{t}^{+}$, and $\Delta_{t}^{-}$, representing the aggregated updates contributed by the clients in $S_{t}^{-}$, as follows:
$\Delta_{t}^{+}:=\frac{1}{n}\sum_{i\in S_{t}^{+}}\Delta_{t}^{i}=\frac{1}{n}\sum_{i\in S_{t}^{+}}|D_{i}|(w_{t}^{i}-w_{t})$ (5)
$\Delta_{t}^{-}:=\frac{1}{n}\sum_{j\in S_{t}^{-}}\Delta_{t}^{j}=\frac{1}{n}\sum_{j\in S_{t}^{-}}|D_{j}|(w_{t}^{j}-w_{t})$ (6)
We also define the forget data (or target data) as the union of the local dataset held by clients that request unlearning, expressed as $D_{u}:=\bigcup_{j\in S_{t}^{-}}D_{j}$.
Figure 2: Overview of PUF operating modes. In PUF-Regular Round, unlearning clients provide their model updates together with other clients. In PUF-Special Round, only unlearning clients share their model updates.
IV-B Federated Unlearning via Negated Pseudo-Gradients
In FedAvg [9], clients receive the weights of the current global model and send back model updates. As discussed in Section III, these updates can be interpreted as pseudo-gradients, also enabling the use of server-side optimizers [21]. We leverage this interpretation for client unlearning in PUF. Specifically, PUF flips the sign of the model update computed by the unlearning client(s), applying it in the opposite direction of the local optimization. Figure 2 provides an illustrative overview our method.
This approach emulates the effect of gradient ascent-based unlearning techniques by pushing the global model parameters away from regions where the loss function on the forgetting data is minimized. A key technical advantage of PUF over directly applying gradient ascent—i.e., maximizing the loss function instead of minimizing it on the forgetting data—is that it avoids the risk of gradient explosion. Since the loss function generally lacks an upper bound, gradient ascent for unlearning can lead to uncontrolled gradient growth, preventing convergence and severely degrading the model’s generalization ability. To address this, methods such as weight projection (e.g., L2-norm ball) and gradient clipping are often employed to constrain parameter updates. However, these techniques require careful and often extensive hyperparameter tuning, including selecting an appropriate L2-norm radius or gradient clipping threshold.
In contrast, our method eliminates the risk of uncontrolled gradient growth by preserving the standard local training procedure on clients while simply negating and rescaling the resulting update (pseudo-gradient) after local training. PUF supports two alternative operating modes:
  • Regular Round with Modified Aggregation (PUF-Regular). In this mode, unlearning is integrated into a regular training round, where both unlearning and remaining clients participate. When one or more clients opt to leave the federation and be forgotten, the round proceeds as usual, except that the aggregated model updates from unlearning clients are negated. The sign flip can be applied either locally at clients, before transmitting back the updates, or during the server-side aggregation. The updates from other clients remain unchanged. The resulting aggregation can be formalized as follows:
    $\Delta_{t}=\eta_{r}\Delta_{t}^{+}-\eta_{u}\Delta_{t}^{-},$ (7)
    with $\eta_{r}$ and $\eta_{u}$ being global learning rates. In particular, $\eta_{u}$ scales the aggregated update from unlearning clients. Considering SGD as a server-side optimizer and $\eta_{s}=1$ (see Eq. 3), the resulting unlearning model is generated as:
    $w_{\bar{u}}=w_{t}+\Delta_{t}$ (8)
    Regular FL training resumes from $w_{\bar{u}}$.
  • Special Unlearning Round with Only Unlearning Clients (PUF-Special). In this mode, a dedicated unlearning round is conducted before resuming regular training. Only the unlearning clients participate, retrieving the latest global model and performing regular training. Assuming SGD as the server-side optimizer with $\eta_{s}=1$, the aggregated update is negated and applied to the global model as follows:
    $w_{\bar{u}}=w_{t}-\eta_{u}\Delta_{t}^{-},$ (9)
    where $\eta_{u}$ is the scaling factor for the aggregated update from unlearning clients. Regular FL training resumes from $w_{\bar{u}}$.
The effectiveness of PUF depends on the magnitude of the unlearning updates, which must be carefully controlled to prevent excessive model drift. As we discuss in Section V, the tuning of the scaling factor $\eta_{u}$ is crucial for balancing unlearning effectiveness while mitigating degradation in the model’s generalization performance.
PUF efficiently removes contributions from unlearning clients while remaining fully compatible with the traditional FedAvg. After the unlearning round, the global model may require a few additional rounds to recover its generalization performance, as observed in all FU mechanisms [13]. Algorithm 1 reports a unified framework for both PUF-Special and PUF-Regular. For readability, at line 14, we assume SGD with $\eta_{s}$=1.0 as the server-side optimizer.
It is worth noting that in both operating modes, clients can transmit model updates to the server, as done in traditional FedAvg, without requiring any modifications or additional overhead. PUF is designed to ensure the unlearning process is task-agnostic and seamless for federated participants, supporting multiple concurrent unlearning requests. Furthermore, as the server performs unlearning, our method also integrates seamlessly with existing security mechanisms.
Algorithm 1 PUF Algorithm. Note that $S_{t}^{+}=\emptyset$ for PUF-Special rounds, while $S_{t}^{+}\neq\emptyset$ and $S_{t}^{-}\neq\emptyset$ for PUF-Regular rounds. $S_{t}^{-}=\emptyset$ for regular training rounds.
1:Global model weights $w$, local epochs $E$, global learning rate $\eta_{s}$, unlearning rate $\eta_{u}$, local learning rate $\eta$, batch size $B$
2:Initialize $w_{0}$
3:for each round $t=0,1,2,\dots$ do
4:     $S_{t}^{+}\leftarrow$ (random set of remaining clients)
5:     $\triangleright$ $S_{t}^{+}=\emptyset$ if PUF-Special
6:     $S_{t}^{-}\leftarrow$ (set of unlearning clients)
7:     $\triangleright$ $S_{t}^{-}=\emptyset$ if no unlearning request
8:     for each client $i\in S_{t}^{+}$ simultaneously do
9:         $\Delta_{t}^{i}\leftarrow$ ClientOpt$(i,w_{t})$
10:     for each client $j\in S_{t}^{-}$ simultaneously do
11:         $\Delta_{t}^{j}\leftarrow$ ClientOpt$(j,w_{t})$
12:     $\Delta_{t}^{+}\leftarrow\frac{1}{n}\sum_{i\in S_{t}^{+}}\Delta_{t}^{i}$
13:     $\Delta_{t}^{-}\leftarrow\frac{1}{n}\sum_{j\in S_{t}^{-}}\Delta_{t}^{j}$
14:     $\Delta_{t}\leftarrow\eta_{r}\Delta_{t}^{+}-\eta_{u}\Delta_{t}^{-}$
15:     $w_{t+1}\leftarrow w_{t}+\Delta_{t}$
16:procedure ClientOpt$(k,w_{t})$
17:     $w\leftarrow w_{t}$
18:     $\mathcal{B}\leftarrow$ (split $\mathcal{D}_{k}$ into batches of size $B$)
19:     for each local epoch $e$ from 1 to $E$ do
20:         for each batch $b\in\mathcal{B}$ do
21:              $w\leftarrow w-\eta\nabla\ell(w;b)$
22:     $\Delta_{t}^{k}\leftarrow w-w_{t}$
23:     return $\Delta_{t}^{k}$ to server
V Experimental Results
V-A Datasets, Models, and Learning Setting
We conducted a comprehensive set of experiments on image classification tasks and image segmentation tasks. For image classification, we used federated versions of the CIFAR-10 and CIFAR-100 datasets [28], which consist of 60,000 32×32 color images—50,000 for training and 10,000 for testing. CIFAR-10 comprises 10 classes, while CIFAR-100 includes 100 classes. The datasets were partitioned to simulate 10 clients, ensuring no overlapping samples among them. The experiments were performed under both Identically and Independently Distributed (IID) and non-IID settings across clients. To create non-IID data, we applied label-skew partitioning based on a distribution determined by Latent Dirichlet Allocation (LDA) [29], using concentration parameters of $\alpha=0.3$ for CIFAR-10 and $\alpha=0.1$ for CIFAR-100. Figure 3 provides a visual depiction of the label distribution among clients. For image classification, we employed a standard ResNet-18 [30] and a visual transformer, namely the MiT-B0 [31]. Before performing unlearning, we train ResNet-18 from scratch for 200 FL rounds while we fine-tune MiT-B0 for 50 FL rounds, starting from a pre-trained checkpoint. If not differently indicated, clients run one local epoch ($E=1$) for standard FedAvg.
For image segmentation, we conducted a set of experiments using the ProstateMRI federated dataset [32] that comprises prostate T2-weighted MRI scans (with segmentation masks) sourced from six data providers, each one treated as a client. Due to variations in imaging protocols across sites, real-world feature heterogeneity naturally arises among clients. The dataset includes 384x384 color images alongside their corresponding segmentation masks.
Table II outlines the settings of our experiments. For each setting, we performed a variable amount of experiments, considering one or more clients to forget. In all experiments, we consider SGD as a server-side optimizer with $\eta_{s}$=1.0, and we set $\eta_{r}$=1.0 for PUF-Regular. If not differently indicated, we set $\eta_{u}$=2.0 for PUF-Special and $\eta_{u}$=20.0 for PUF-Regular (Section V-E discusses the tuning of our methods). Specific detail about per-setting hyper-parameters are reported in the following paragraphs.
(a) CIFAR-100 (IID).
(b) CIFAR-10 (Non-IID, $\alpha=0.3$).
(c) CIFAR-100 (Non-IID, $\alpha=0.1$).
[…]
Model Dataset Clients Data Heterogeneity Pre-trained $E$ Task # Unlearning Clients ResNet-18 CIFAR-10 10 LDA ($\alpha$=0.3) $\times$ 1 $\mathcal{C}$ Up to one ResNet-18 CIFAR-100 10 LDA ($\alpha$=0.1) or IID $\times$ 1 or 10 $\mathcal{C}$ Up to two MiT-B0 CIFAR-100 10 LDA ($\alpha$=0.1) 1 $\mathcal{C}$ Up to two ResUNET ProstateMRI 10 Real-world feature skew $\times$ 1 $\mathcal{S}$ Up to one
TABLE II: Settings of reported results. $\mathcal{C}$=classification task, $\mathcal{S}$=segmentation task. $E$=local epochs during FedAvg. # Unlearning Clients reports if the experiments considers a single client or multiple clients to unlearn.
ResNet-18 on CIFAR-10/CIFAR-100. We used a standard ResNet-18 [30] and employed Group Normalization layers, similar to other works with similar settings (e.g., [33]). We used SGD as a local optimizer with a learning rate set to 0.1, with a round-wise exponential decay of 0.998, 1 or 10 local epochs ($E=1$ or $E=10$), local batch size of 32. We pre-processed the training images with random crop, horizontal flip and normalization layers. During unlearning routines, we only apply normalization.
MiT-B0 on CIFAR-100. We used a visual transformer, i.e., MiT-B0 [31], with approximately 3.6M parameters, initialized from a pre-trained model checkpoint trained on ImageNet-1k (69.27% accuracy on test data). We adapted the one-layer classification head to this task, initializing such a layer from scratch. We employed the AdamW optimizer with a client learning rate of 3e-4, with a round-wise exponential decay of 0.998, 1 local epoch, local batch size of 32, and weight decay regularization of 1e-3. The images are resized to a resolution of 224x224.
Res-UNet on ProstateMRI. We utilized a vanilla Res-UNet architecture [34], similar to the one used in [32], with approximately 7.6M parameters. In line with [35], we trained the network using a combination of standard cross-entropy and Dice loss [36], and we conducted regular training for 500 rounds, before applying unlearning. We used Adam [37] as local optimizer with a client learning rate of 1e-4, a local batch size of 16, and a local weight decay of 1e-4.
V-B Baselines
We compared PUF against two state-of-the-art baseline methods, which are PGA [18], and NoT [19]. These methods were selected because, similarly to PUF, they do not depend on impractical assumptions, such as access to supplementary data or historical client updates (as we detail in Section II). In fact, requiring a complete history of client updates would necessitate linking each client to their previous submissions, thereby compromising the privacy-centric design of FL. Similarly, assuming the availability of useful public data on the server side to restore the performance of the unlearned model is a strong assumption that is often unrealistic in practice.
Since PGA requires full client participation, we ensured that the original model was trained with full client participation to guarantee a fair comparison. In some experiments, we included a Natural baseline, which bypasses any unlearning procedure before the recovery phase. This allows us to assess whether the influence of the unlearning client would naturally vanish over time without intervention.
V-C Metrics
We evaluated the methods based on two key aspects: their efficiency in recovering performance and their effectiveness in achieving unlearning. When evaluating the effectiveness of unlearning, the primary criterion is minimizing the discrepancy with the metrics of the gold-standard retrained model. This means that neither higher nor lower metric values are inherently preferable; rather, the objective is to reduce the absolute difference between the metrics of the approximate unlearning method and those of the retrained model. We use $\Delta$ as a prefix to denote the absolute difference from the retrained model’s value for a specific metric (e.g., $\Delta$ Test Accuracy represents the absolute difference in Test Accuracy).
Test Accuracy ($\Delta$ Test Acc.). We evaluated the test accuracy at two critical stages: immediately after the unlearning routine (before the recovery phase) and after the recovery phase concludes. We used the standard test set provided with the dataset, which is IID and contains data that have never been accessed or observed by any client during training.
Rounds to Recover Performance (Recovery Rounds). We measured the number of FL rounds required by each evaluated method to recover or exceed the test accuracy of the corresponding retrained model. We refer to such a recovery period as recovery phase. Lower values indicate faster recovery, which is preferable.
Communication Efficiency (CE) over retraining. We also report the communication reduction relative to producing a retrained model. Since none of the baselines we consider introduce additional per-round communication overhead compared to FedAvg, we compute CE as the number of rounds needed to learn a retrained model over the recovery rounds required by the specific FU algorithm.
Setting Algorithm Rounds ($\downarrow$) CE ($\uparrow$) Test Acc. Forget Acc. MIA [38] MIA [39] CIFAR-100, IID, $E$=1, ResNet-18 Original 59.91±0.0 79.82±0.7 78.23±0.6 71.42 ±0.6 Retrain 58.32±0.5 58.01±0.7 55.62±0.7 48.86 ±0.6 PGA [18] 12.1 ±4.2 16.7$\times$ 59.12±0.6 72.81 (14.81±1.0) 71.83 (16.15±1.8) 63.02 (14.12±1.6) NoT [19] 6.51 ±1.1 30.8$\times$ 58.94 ±0.4 70.51 (12.61±0.8) 69.31 (13.72±1.9) 63.34 (14.47±0.8) PUF-Special 4.1±1.0 48.8$\times$ 58.82 ±0.3 62.84 (4.86±1.08) 61.49 (5.88±1.42) 55.82 (6.98±1.13) PUF-Regular 5.2±0.63 38.5$\times$ 58.78±0.4 62.36 (4.39±1.01) 60.83 (5.23±0.92) 55.52 (6.68±1.07) CIFAR-100, Non-IID, $E$=1, ResNet-18 Original 53.81±0.0 62.93±6.9 75.44±7.0 61.04 ±7.8 Retrain 51.02±1.4 33.52±4.5 44.02±5.5 32.05 ±5.2 PGA [18] 7.0±5.1 28.6$\times$ 51.42 ±0.6 37.43 (4.32±4.1) 49.21 (6.44±4.90) 36.62 (5.33±5.3) NoT [19] 12.8±6.8 15.6$\times$ 51.31 ±0.2 43.31 (9.9±4.4) 55.75 (11.7±5.7) 44.32 (12.3±5.5) PUF-Special 7.3±4.14 27.4$\times$ 51.49±0.4 35.04 (2.42±2.04) 46.74 (3.66±2.49) 34.34 (2.72±2.22) PUF-Regular 8.0±4.16 25.0$\times$ 51.28±1.46 33.74 (2.37±1.76) 45.51 (3.07±2.42) 33.09 (2.14±2.08) CIFAR-100, Non-IID, $E$=10, ResNet-18 Original 50.32±0.0 60.81±6.63 73.42±6.6 61.31 ±7.5 Retrain 48.02±0.8 31.61±4.5 42.12±5.2 31.24 ±4.9 PGA [18] 4.4±3.8 45.5$\times$ 49.02±1.0 37.34 (6.27±8.1) 49.65 (7.7±9.6) 37.72 (6.6±8.2) NoT [19] 7.7±1.49 26.0$\times$ 48.71±0.7 42.13 (10.5±3.3) 55.72 (13.6±3.3) 45.01 (13.8±4.2) PUF-Special 5.4±1.26 37.3$\times$ 48.36±0.75 32.75 (3.11±2.22) 44.71 (3.71±2.6) 33.92 (3.99±2.28) PUF-Regular 6.3±1.25 31.8$\times$ 48.53 ±0.88 32.48 (1.88±1.6) 43.97 (2.03±1.88) 33.17 (2.44±1.8) CIFAR-10, Non-IID, $E$=1, ResNet-18 Original 83.72 ±0.02 88.64 ±3.93 84.41 ±5.42 81.41 ±6.33 Retrained 83.52 ±1.64 81.11 ±8.04 73.16 ±13.71 72.52 ±10.82 PGA [18] 11.5±5.6 17.4$\times$ 84.01 ±1.0 84.31 (3.20±4.9) 78.54 (5.40±7.7) 77.02 (4.51±6.0) NoT [19] 16.9±8.6 11.8$\times$ 83.81 ±1.2 84.31 (3.10±3.0) 78.04 (4.91±4.5) 75.33 (2.90±3.4) PUF-Special 10.4±6.0 19.2$\times$ 83.81 ±0.2 82.32 (2.01±1.96) 73.33 (3.21±2.72) 73.15 (2.54±2.16) PUF-Regular 13.1±12.1 15.3$\times$ 83.72±1.6 82.22 (1.90±1.5) 73.01 (3.6±3.4) 73.03 (2.52±2.1) CIFAR-100, Non-IID, $E$=1, MiT-B0 Original 75.03 ±0.00 84.25 ±5.63 77.03 ±8.57 74.03 ±6.43 Retrained 73.30 ±0.78 57.80 ±5.29 48.59 ±4.96 44.59 ±3.88 PGA [18] 6.9 ±4.28 7.2$\times$ 73.60 ±0.36 62.15 (4.35±3.28) 53.57 (4.98±3.67) 50.50 (6.91±3.21) NoT [19] 15.5±10.17 3.2$\times$ 73.44±0.74 67.85 (10.05±4.07) 59.42 (10.07±6.59) 53.99 (9.40±5.39) PUF-Special 13.8±5.09 3.6$\times$ 73.55±0.7 60.01 (2.54±1.37) 51.04 (2.46±1.24) 46.71 (2.11±1.11) PUF-Regular 17.0±8.27 2.9$\times$ 73.44±0.76 59.52 (2.00±1.17) 50.61 (1.97±1.18) 47.29 (2.11±1.15)
TABLE III: Performance of PUF and other baselines across different settings, including data distribution and the number of local epochs, with ResNet-18 or MiT-B0. Communication Efficiency (CE column). Baseline results are expressed as mean metric value (mean $\Delta$$\pm$ standard deviation), with $\Delta$ in parentheses representing the average absolute difference from the retrained model, highlighted in blue for better readability. Lower $\Delta$ corresponds to better unlearning performances. Note that, the baselines’ Test Acc. is, by definition, higher than that of the retrained model, as we consider the recovery phase complete once Test Accuracy surpasses that of the retrained model.
Forget Accuracy ($\Delta$ Forget Acc.). We evaluated the unlearned model’s accuracy on client $u$’s train data and computed the absolute difference relative to the retrained model’s performance. This metric provides insight into how closely the unlearning process approximates the desired outcome achieved through retraining.
MIAs on Forget Data ($\Delta$ MIA). The MIA aims to infer whether specific samples were part of the training data for the attacked model. Its success rate quantifies the fraction of target data, defined as ($D_{u}:=\bigcup_{j\in S_{t}^{-}}D_{j}$), correctly identified as belonging to the training set. A lower MIA success rate indicates reduced information leakage about $D_{u}$ from the attacked model. To implement MIAs, we employed two established approaches: (i) a confidence-based MIA predictor [38], and (ii) a loss-based MIA predictor [39].
  1. 1.
    Confidence-based MIA predictor [38]: It consists of a training phase using a balanced dataset of seen and unseen data. The retain dataset serves as the seen data, while the standard test dataset is used as the unseen data. An MIA predictor is then trained to classify whether the target model’s output corresponds to seen or unseen data based on prediction confidence.
  2. 2.
    Loss-based MIA predictor [39]: It assumes the attacker can access the target model’s average training loss. Samples are classified as members of the training set if their loss falls below this average; otherwise, they are labeled as non-members. For our federated implementation, we assume access to the global mean training loss.
(a) CIFAR-100, IID, ResNet-18.
(b) CIFAR-100, Non-IID, ResNet-18.
(c) CIFAR-100, non-IID, MiT-B0.
Figure 4: Performance comparison of PUF with baselines for different experimental settings indicated in subcaption as a triple dataset, data distribution, model architecture. A smaller polygon represents better unlearning performances. For our method, the charts only visualize the performance of PUF-Special for better visualization. As reported in Table III with the complete results, PUF-Regular performs very similarly to PUF-Special.
(a) CIFAR-100, IID, ResNet-18.
(b) CIFAR-100, Non-IID, ResNet-18.
(c) CIFAR-100, Non-IID, MiT-B0.
Figure 5: Evolution of generalization ability (test accuracy) and forgetting effectiveness (forget accuracy) during the recovery phase across three different settings. Each pair of images presents Test Accuracy (Left) and Forget Accuracy (Right) for a representative client in a specific setting indicated in subcaption as a triple dataset, data distribution, model architecture. For our method, the charts only report the performance of PUF-Special for better visualization.
V-D Main Results
Table III presents the performance of PUF best configurations (hyper-parameter tuning discussed in Sec. V-E) and other baseline methods on different datasets (federated CIFAR-100 and CIFAR-10 datasets), data distributions, and model architectures (ResNet-18, MiT-B0). When not differently stated, the unlearning task considers a single client, and the results are averaged over ten independent experiments, each involving a different client designated as the unlearning client. For results on multiple unlearning (two unlearning clients), the results are averaged over five independent experiments, selecting the set of unlearning clients randomly.
Overall Performance. Figure 4 provides a visual comparison of the performance of PUF-Special against other baselines, where a polygon with a smaller area indicates better overall performance. Forgetting effectiveness is assessed using three metrics: the average absolute gap ($\Delta$) in Forget Accuracy and two variations of MIAs. The efficiency of each FU method is represented by the number of Recovery Rounds required. Additionally, the chart evaluates the FU method’s ability to preserve the global model’s generalization. This is measured by the degradation in Test Accuracy introduced by the unlearning process, quantified as the difference in Test Accuracy between the retrained model and the unlearned model immediately after unlearning is applied — before recovery begins (ideally, at round 0 of recovery).
Overall, PUF demonstrates (1) superior forgetting effectiveness compared to other baselines, and (2) comparable or superior efficiency, except in the CIFAR-100, non-IID, MiT-B0 setting (see Figure 4.c and the 4th row of Table III), where PUF requires a longer recovery period than PGA but achieves significantly better forgetting effectiveness.
Performance Consistency During Recovery. Figure 5 illustrates the evolution of Test Accuracy and Forget Accuracy throughout the recovery phase, beginning at recovery round 0, which corresponds to the application of the unlearning routine (either PGA, NoT, or our proposed solution, PUF). We do not report the evolution of the considered MIAs, as they closely follow the trends of Forget Accuracy.
The figure presents results for three different settings, as indicated in the subcaptions. For reference, we also include values for both the Original Model and the Retrained Model. We emphasize the following key aspects for the full and easy understanding of the reported charts:
  1. 1.
    The recovery phase concludes when the unlearned model surpasses the Test Accuracy of the retrained model for the first time.
  2. 2.
    A more effective forgetting method should minimize the gap in Forget Accuracy at the end of the recovery phase.
  3. 3.
    A better forgetting method should also reduce the discrepancy with the retrained model’s Test Accuracy during the recovery phase, ensuring higher usability of the unlearned model, which serves as the new FL global model for inference during these rounds.
Thus, Figure 5 provides valuable insights into the dynamics of these metrics during the recovery phase. The analysis highlights that PUF produces a superior unlearned model, significantly narrowing the gap with the retrained model earlier in the recovery phase compared to the other baselines. Additionally, the figure clearly shows that the other baselines fail to achieve accurate forgetting by the end of the recovery phase, as evidenced by their larger gaps in Forget Accuracy.
Per-round Computational and Communication Cost. PUF’s unlearning round has the same computational and communication requirements as a regular FedAvg training round, as both PUF-Special and PUF-Regular achieve effectiveness using the same number of local epochs as standard training.
On the other hand, PGA requires substantially more local epochs to generate the unlearned model—up to 5$\times$—to achieve satisfactory unlearning performance, as reported in its original studies. NoT does not require any client-side computation to produce the unlearning model but exhibits less efficiency in terms of longer recovery phase.
Cumulative Computational and Communication Cost. NoT requires significantly more recovery rounds to restore the performance of the unlearned model, resulting in lower cumulative computational and communication efficiency compared to PUF (see the CE column in tabular results). In the CIFAR-100, Non-IID, MiT-B0 setting, PGA demonstrates greater efficiency than PUF, requiring fewer recovery rounds and incurring lower cumulative costs. However, it produces a less effective global model in terms of forgetting metrics, as already highlighted above in this discussion.
Seamless Integration into FedAvg. PUF-Regular can be effortlessly incorporated into regular FedAvg rounds by reversing the direction of the unlearning clients’ (aggregated) updates. Unlike methods such as PGA, or PUF-Special, PUF-Regular does not alter the regular training process or require a dedicated unlearning round to remove clients’ contributions.
Setting Algorithm Rounds ($\downarrow$) CE ($\uparrow$) Test Acc. Forget Acc. MIA [38] MIA [39] Multiple Unlearning, CIFAR-100, Non-IID, $E$=1, ResNet-18 Original 53.81±0.0 63.06±6.83 75.95±7.55 61.09±8.32 Retrain 47.32±2.13 30.09±3.29 39.66±3.90 27.89±3.12 NoT [19] 6.6±1.82 30.0$\times$ 48.06±2.24 39.16 (9.07 ±2.83) 50.51 (11.26 ±3.39) 41.8 (13.9 ±4.52) PUF-Special 3.2±1.1 62.5$\times$ 48.53±2.14 34.12 (4.04 ±3.69) 46.09 (6.64 ±3.46) 33.5 (5.61 ±4.21) PUF-Regular 8.0±2.83 25.0$\times$ 47.85±2.3 31.24 (2.87 ±1.86) 42.85 (3.89 ±2.94) 31.35 (3.78 ±3.13) Multiple Unlearning, CIFAR-100, Non-IID, $E$=1, MiT-B0 Original 75.03±0.0 84.60±5.25 76.84±8.95 73.8±1.40 Retrain 70.86±1.32 54.30±3.49 45.36±3.64 43.5±1.61 NoT [19] 3.2±3.03 15.6$\times$ 71.38±0.94 69.89 (15.58 ±6.29) 62.52 (16.55 ±8.89) 58.95 (15.45 ±7.50) PUF-Special 6.0±1.22 8.3$\times$ 71.28±1.44 60.58 (6.28 ±2.08) 52.97 (7.25 ±2.71) 50.50 (7.01 ±1.84) PUF-Regular 16.0±3.04 3.1$\times$ 70.87 ±1.49 56.57 (2.26 ±1.07) 48.67 (2.55 ±2.32) 45.62 (2.10 ±1.12)
TABLE IV: Performance of PUF and other baselines for multiple unlearning (unlearning two clients simultaneously). The column meaning is the same as Table III. Lower $\Delta$ corresponds to better unlearning performances.
Multiple Unlearning. Table IV presents the performance on federated CIFAR-100 (non-IID) using a ResNet-18 or an MiT-B0 architecture, where a pair of randomly chosen clients was designated as unlearning clients (averaged over five independent experiments). We do not include PGA as a baseline because it is not directly extensible to unlearning multiple tasks. This limitation arises from its reliance on a reference model, which is obtained by removing the last stored model update of the target client. As a result, the mutual dependency of unlearning updates prevents its straightforward adaptation to multiple unlearning requests.
As shown in Table IV, the most effective unlearning of multiple clients simultaneously is achieved by either PUF-Special or PUF-Regular. While NoT benefits from a relatively short recovery phase, it demonstrates significantly weaker forgetting performance, as indicated by larger gaps in Forget Accuracy and MIAs. Notably, even when handling multiple unlearning requests, PUF-Regular, like NoT, operates without requiring any modifications to the standard FedAvg protocol.
Algorithm Rounds ($\downarrow$) CE ($\uparrow$) Test Acc. Forget Acc. Forget Loss Original 86.54±0.0 88.94±4.58 0.08±0.05 Retrained 82.69±1.68 73.29±7.23 0.22±0.10 NoT [19] 58.6±22.74 8.5$\times$ 83.43 ±0.74 77.37 (3.91±2.42) 0.17 (0.06±0.02) PUF-Special 41.8±11.95 12.0×12.0$\times$12.0 × 83.06±0.37 76.71 (3.38±2.27) 0.17 (0.05±0.02)
TABLE V: Performance of PUF and other baselines on ProstateMRI dataset. Values in parenthesis represent the mean absolute difference with the retrained model, where a lower value is better.
Task Agnosticism. As highlighted in Sections II and IV, the design of PUF ensures that the unlearning phase remains task-agnostic, as both the unlearning and remaining clients follow the standard training procedure without any modifications to their local learning routines. Table V reports the results for the ProstateMRI experiments, which focus on a segmentation task rather than classification. PUF outperforms all baselines in terms of both recovery efficiency (i.e., shorter recovery phase) and forgetting efficacy. As discussed in Section V-E, PUF can be further tuned for improved forgetting performance by increasing the scaling factor $\eta_{u}$, albeit at the cost of a longer recovery phase. Notably, NoT does not offer this tuning flexibility.
(a) CIFAR-100, Non-IID, ResNet18.
(b) CIFAR-10, Non-IID, ResNet18.
(c) CIFAR-100, Non-IID, MiT-B0.
Figure 6: Performance of PUF with varying hyper-parameters ($\eta_{u}$, $E_{u}$) compared to related methods. When $E_{u}$ is omitted, it is set to 1. X-axis: number of recovery rounds required to match the retrained model’s test accuracy. Y-axis: gap in forget accuracy compared to the retrained model. Points closer to the origin indicate better performance. The experimental setting is indicated in subcaption as a triple dataset, data distribution, model architecture. The Natural baseline reports results for a naive strategy of fine tuning the global model without the participation of the unlearning client, no explicit unlearning is performed.
Figure 7: Performance of PUF with varying hyper-parameter $\eta_{u}$ compared to related methods for the ProstateMRI setting.
V-E Hyper-parameter Tuning of PUF
Figure 6 compares the unlearning performance ($\Delta$ Forget Accuracy on the Y-axis, recovery rounds on the X-axis) of various configurations of PUF-Special and PUF-Regular with the best configurations of NoT and PGA, and the Natural baseline. The analysis also includes configurations where PUF employs more local epochs (denoted as $E_{u}$ in Figure 6; when omitted, $E_{u}=1$). Points closer to the origin of the axes indicate better performance. For this analysis, we focused exclusively on the $\Delta$ Forget Accuracy metric, as the MIAs metrics align closely with this metric (as emerges from tabular results). We highlight in red the configurations we found to be the best and used them in the Experimental Results section.
Notably, increasing local epochs ($E_{u}>1$) for PUF reduces $\Delta$ Forget Accuracy but requires more recovery rounds compared to $E_{u}=1$ with the same $\eta_{u}$. Interestingly, increasing the scaling factor $\eta_{u}$ while keeping $E_{u}=1$ achieves similar improvements without additional local computational overhead for unlearning clients, making it more practical in resource-constrained settings. Both PUF-Regular and PUF-Special achieve high effectiveness by simply tuning the scaling factor ($\eta_{u}$) while maintaining $E_{u}=1$. As shown in Figure 6, PUF demonstrates low sensitivity to the $\eta_{u}$ hyper-parameter, resulting in a cluster of well-performing configurations.
In contrast, we observe that PGA requires tuning multiple hyper-parameters – local learning rate, local epochs, an early stopping threshold, and gradient-clipping threshold – while also assuming full client participation and stateful clients. On the other hand, NoT does not require specific hyperparameter tuning but fails to achieve significant forgetting. As shown in Figure 6.a, NoT performs only slightly better than the Natural baselines in both recovery efficiency and forgetting effectiveness. This highlights PUF’s trade-off between simplicity, ease of hyperparameter optimization, and unlearning performance compared to other baselines.
Figure 7 shows the hyper-parameter tuning of PUF for the ProstateRMI dataset, on segmentation data and real-world feature heterogeneity. Also in this setting, PUF exhibits low sensitivity to its main hyper-parameter.
VI Conclusive Remarks
In this paper, we introduced PUF, a novel FU method that leverages negated pseudo-gradients to enable clients to exercise their right to be forgotten. Aligned with the design principles of FL, PUF does not require storing any processed client information and relies only on ephemeral updates, even during unlearning. It integrates seamlessly with standard FedAvg, by requiring no significant modifications or additional computational/communication overhead. As a result, PUF is both task-agnostic and capable of handling multiple unlearning requests concurrently, which are two critical aspects that most existing literature does not consider. Experimental results on two classification datasets and a segmentation task, across varying degrees of data heterogeneity and different model architectures, are reported to show the general applicability of the proposed approach: PUF has clearly demonstrated to be more efficient in recovering expected performance and more accurate in inducing forgetting of requested data compared to state-of-the-art baselines.